OpenStudy (anonymous):

A university bookstore recently sold a wirebound graph-paper notebook for $2.55, and a college-ruled notebook for $3.22. At the start of spring semester, a combination of 50 of these notebooks were sold for a total of $157.65. How many of each type were sold?

OpenStudy (anonymous):

Doyou know how to solve systems of linear equations?

OpenStudy (anonymous):

using simultaneous eq's let x = graph book let y = ruled book since 50 were sold in total---> x + y = 50 and since the total earnings were 157.65 ---> 2.55x +3.22y = 157.65 multiply the top eq. by 3.22 so that the x terms can be eliminated... therefore you get : 3.22x + 3.22y = 161 - 2.55x + 3.22y = 157.65 after subtraction you get 0.67x = 3.35 therefore x = 5 substitute the value back into original eq : x + y = 50 so y = 45